A photon is emitted when a hydrogen atom undergoes a transition from the n=5 state to the n=3 state. Calculate (a) the energy in eV (b) the wavelength (c) the frequency of the emitted photon
1
Expert's answer
2020-10-15T10:50:45-0400

Solution

a) energy in case of Hydogen atom is by

"E=13.6 (\frac {1}{n_2^2}-\frac {1}{n_1^2}) eV"

Here"\space n_1=5\space \space n_2=3" so energy is given by

"E=13.6 (\frac {1}{9}-\frac {1}{25}) eV"

"E=0.98eV"

b) wavelength to corresponding energy is given by

"E=\frac{hc}{\lambda}"

"\lambda=\frac{hc}{E}"

By putting the value of h, c, E

Wavelength becomes

"\lambda=12.63\times10^{-7}m"

c) now frequency is given by

"\upsilon=\frac{E}{h}"

Now putting the value of E and h

Frequency is given by

"\upsilon=0.24\times10^{15} Hz"